# Coordinated UAV Beamforming and Control for Directional Jamming and Nulling

### v1

@@ Abstract
Efficient mobile jamming against eavesdroppers in wireless networks necessitates accurate coordination between mobility and antenna beamforming. We study the coordinated beamforming and control problem for a UAV that carries two omnidirectional antennas, and which uses them to jam an eavesdropper while leaving a friendly client unaffected. The UAV can shape its jamming beampattern by controlling its position, its antennas’ orientation, and the phases of the antennas’ interference signals. We derive a closed-form expression for the antennas’ phases that guarantees zero jamming impact on the client. In addition, we determine the antennas’ orientation and the UAV’s position that maximizes jamming impact on the eavesdropper through an optimal control problem, optimizing the orientation pointwise and the position through the UAV’s control input. Simulations show how this coordinated beamforming and control scheme enables directional GPS denial while guaranteeing zero interference towards a friendly direction.
@@ I Introduction
Two factors largely determine the effectiveness of a jamming signal: the distance to the target and the shape of the antenna’s radiation pattern. Because signal strength decays with distance, staying close to the target ensures sufficient interference power. At the same time, the antenna’s beampattern, i.e., the way its transmitted energy is distributed over different directions, dictates where that power is delivered. A well-shaped beam can focus interference on an eavesdropper, whereas a poorly shaped one may also radiate energy toward friendly receivers. Figure 1 illustrates this fact, showing how beamshaping can direct interference selectively at an adversary while avoiding spread towards unintended directions.
Beamshaping for directional jamming has been studied extensively in wireless communication and antenna theory; however, its combination with dynamics is relatively limited. For example, [7, 8, 9, 10, 11, 12] studied antenna placement and beampattern design for jamming, but not the case where the antenna carrier is a mobile UAV. A related strand incorporates UAV mobility by optimizing trajectories [13, 14, 15, 16], but typically in slot-based formulations that ignore vehicle dynamics and actuator limits, and can generate beampatterns that radiate interference toward friendly bases. These approaches highlight the value of beamshaping and UAV mobility, but leave open the question of how to integrate beamforming with dynamics, control, and nulling constraints in a unified framework.
@@
Research in control systems, in contrast, has emphasized the role of mobility in performing or defending against jamming attacks; however, the underlying communication model is often simplified and separated from mobility. For example, [17, 18, 19, 20, 21] presented strategies for planning or agent placement to perform jamming, but assume disk models [^1] Disk model is the assumption that jamming becomes uniformly impactful within a radius from the target, and is inefficient outside that radius., single omnidirectional antennas, or fixed beampatterns. These assumptions are often realistic, but restrict capabilities to cases where the target is far from friendly bases or in line of sight; otherwise, jamming the target also jams the friendly base. Other studies have modeled jamming at the control system level [22, 23, 24, 25, 26, 27, 28, 29], but do not account for the communication model that relates spatial geometry to jamming impact. This gap motivates our study of coordinated beamforming and control for directional jamming and nulling.
We consider a unified framework that integrates antenna beamforming with dynamics to achieve directional jamming with guaranteed nulls toward friendly clients. Specifically, we study a jammer UAV that carries two omnidirectional antennas, and which controls their phases, orientation, as well as its position to i) ensure zero jamming power reaches a friendly client, and to ii) maximize jamming power radiated towards an eavesdropper. The use of two antennas here enables the formation of smooth beampatterns that are robust to variations in the eavesdropper’s direction, as shown in Figure 1.
@@
Notation: $\mathds{1}_{x\in A}(x)$ is the indicator function that is equal to $1$ when $x\in A$, and 0 otherwise. $\left\lVert\cdot\right\rVert$ denotes the Euclidean norm, and $\left\lVert\cdot\right\rVert_{\infty}$ the infinity norm. The function $\textrm{atan2}(y,x):\mathbb{R}\times\mathbb{R}\rightarrow[0,2\pi]$ denotes the two-argument arctangent. We use $I$ to denote an identity matrix of appropriate order, and $\textrm{diag}(c_{1},\ldots,c_{n})$ to denote a diagonal matrix with entries $c_{1},\ldots,c_{n}$. For a vector $x$, $x_{i}$ denotes its $i$-th entry.
@@ II Problem Formulation
We consider a spatial setting where a client wants to receive a wireless communication signal with wavelength $\lambda>0$, while an eavesdropper aims to intercept the same signal. We denote the position of the client in this setting as $p_{c}=[x_{c}~y_{c}]^{\textrm{T}}\in\mathbb{R}^{2}$, and the position of the eavesdropper as $p_{e}=[x_{e}~y_{e}]^{\textrm{T}}\in\mathbb{R}^{2}$.
@@ II-A Spatial Setup for Coordinated Jamming
To prevent the eavesdropper from intercepting the communication, a UAV carries two omnidirectional antennas in order to transmit electromagnetic interference, i.e., to jam the eavesdropper. However, for the jamming to be effective, the UAV must appropriately control the phases of its interference signals as well as its position. Specifically, it must control them to ensure that the interference i) arrives as powerful as possible at the eavesdropper, hence maximizing jamming impact; and ii) cancels out at the client, hence leaving the client completely unaffected.
We set the distance between the two antennas to be fixed and equal to $D>0$, so that they form a basic two-element antenna array. Given this, if we denote the position of the UAV as $p_{g}=[x_{g}~y_{g}]^{\textrm{T}}\in\mathbb{R}^{2}$, then we can express the positions $p_{1},p_{2}\in\mathbb{R}^{2}$ of each of the antennas as
@@
where $\theta_{g}$ denotes the angle of the second antenna relative to the reference frame centered at $p_{g}$. We provide a geometric illustration of this setup in Figure 2.
@@
Instead of employing a UAV carrying two omnidirectional antennas, one could also employ two UAVs carrying one antenna each. In this case, however, there may be limitations on the physical sizes of the UAVs depending on the chosen distance $D$ between them.
Figure 2: The considered setup. A UAV positioned at $p_{g}=[x_{g}~y_{g}]^{\textrm{T}}$ carries two omnidirectional antennas. This UAV must control its position $p_{g}$, the angle of the two antennas $\theta_{g}$, as well as the antennas’ phases in order to maximally jam the eavesdropper positioned at $p_{e}=[x_{e}~y_{e}]^{\textrm{T}}$. At the same time, a client positioned at $p_{c}=[x_{c}~y_{c}]^{\textrm{T}}$ should remain unaffected.
@@ II-B Jamming Beampatterns and Power
The jamming beampattern generated by the two antennas reflects the alignment of their phases, which should ideally combine constructively at the eavesdropper $p_{e}$ to maximize jamming effectiveness, and destructively at the client $p_{c}$ to leave it unaffected. At any point $p\in\mathbb{R}^{2}$, this beampattern takes the expression
@@
where $\phi_{1},~\phi_{2}\in\mathbb{R}$ are the phases of the antennas’ interference signals, $d_{1}(p)=\left\lVert p-p_{1}\right\rVert,~d_{2}(p)=\left\lVert p-p_{2}\right\rVert$ are the distances of each of the antennas from $p$, and $k=\frac{2\pi}{\lambda}$.
@@ II-C Coordinated Beamforming and Control for Maximal Jamming
In the case of an immobile UAV, the jamming power at the eavesdropper’s location $p_{e}$ (as given by (3)) can be maximized using standard optimization tools by statically selecting the phases $\phi_{1}$, $\phi_{2}$, and the angle $\theta_{g}$. However, hereron we consider a mobile UAV, where jamming power can be further enhanced by dynamically controlling the UAV position $p_{g}=p_{g}(t)$ over time $t\geq 0$. In this setting, the jamming power (3) becomes time-varying:
@@
and so do the values of the phases $\phi_{1}(t),\phi_{2}(t)$ and the angle $\theta_{g}(t)$, allowing for greater jamming flexibility. Controlling the UAV position $p_{g}(t)$ while concurrently selecting the antennas’ communication parameters $\theta_{g}(t),\phi_{1}(t),\phi_{2}(t)$ is the purpose of the remainder of the paper.
@@
$\quad\qquad\left\lVert u_{g}(t)\right\rVert_{\infty}\leq\bar{u},~\forall t\in[0,t_{f}].$
@@
While the free-space path losses term $L$ indicate that smaller distances lead to higher jamming impact, it is evident from (4) that such impact also depends on the beamforming function $B$. This means that the array formed by the UAV’s antennas should reposition not only to get closer to the eavesdropper, but also arrive from the proper direction and with the appropriate orientation $\theta_{g}$. For this reason, the position $p_{g}$ of the UAV is not penalized explicitly in (5) but, instead, its desired values are informed through the jamming impact power $P$.
@@ III-A Phase Control for Nulling at the Client
guarantees the client remains unaffected by the interference that the antenna array transmits. Subsequently, we may choose the phase $\phi_{1}(t)$ arbitrarily: since it represents just an offset on the phase of the beamformed signal on the eavesdropper, it has no effect on the magnitude of the beampattern (1). However, in case the jamming bandwidth is small, we can adjust $\phi_{1}(t)$ to offset for the Doppler shift $f_{D}$ on the jammer according to
@@ III-C Position Control for Maximal Jamming Impact on the Eavesdropper
Having chosen the jamming signal phases $\phi_{1}(t),~\phi_{2}(t)$ and the array orientation $\theta_{g}(t)$, the final step is to select the control input $u_{g}(t)$ that repositions the UAV for maximal jamming impact. To this end, plugging the expression of the optimal antenna orientation (10) in (9), we obtain the optimal beampattern value:
@@
$\begin{split}{\xi}_{p}(t_{f})&=-a_{f}\gamma(t_{f})\Bigg{(}\frac{B_{p}(p_{g}(t_{f}))\cdot\mathds{1}_{|\mu(t_{f})|<\frac{\pi}{2kD}}}{B^{\star}(\theta_{e}(t);p_{g}(t_{f}))}\\ &\qquad\qquad\qquad\quad{+}\frac{L_{p}(p_{g}(t_{f}))}{L(\left\lVert p-p_{g}(t_{f})\right\rVert)}\Bigg{)},\\ {\xi}_{v}(t_{f})&=Q_{f}v_{g}(t_{f}).\end{split}$
@@ IV-A Directional Jamming Against a Static Eavesdropper
To jam the eavesdropper and prevent it from receiving the wireless signal, we deploy a jammer UAV that carries two omnidirectional antennas separated by $D=\frac{\lambda}{2}=9.52~\textrm{cm}$ and with power $P_{0}=0.6~\textrm{W}$ each, as shown in Figure 2. This UAV is initially positioned at $p_{g}(0)=[0~0]^{\textrm{T}}~\textrm{m}$ and with velocity $v_{g}(0)=[0~0]^{\textrm{T}}~\textrm{m/s}$. This position, however, is suboptimal: the UAV is far from the eavesdropper, and both the client and the eavesdropper lie in the same direction in the far field. As a result, the UAV cannot effectively jam the eavesdropper without also interfering with the client. Therefore, to reposition the UAV for maximum jamming effectiveness, we solve the optimal control problem (14) to determine its control input $u_{g}$, and we select the communication parameters $\phi_{1},~\phi_{2},~\theta_{g}$ according to (6), (7) and (10). The actuation threshold for the UAV is $\bar{u}=2~\textrm{m/s}^{2}$.
@@
and, accordingly,
$\sigma^{\prime}(x)=\begin{cases}0,&x<-100,\\ 1,&x\in[-100,-70],\\ 0,&x>-70.\end{cases}$
This choice of $\sigma$ motivates the optimal $u_{g}^{\star}$ to generate trajectories that degrade or deny service at the eavesdropper. At the same time, it avoids giving excessive rewards for UAV paths that pass too close to the eavesdropper and yield jamming powers above $70~\textrm{dBm}$ (as the DoS outcome remains unchanged). Moreover, it preserves numerical stability by zeroing the flow equations in Theorem 2 whenever either $B^{\star}$ or $\left\lVert p-p_{g}\right\rVert$ is close to zero.
@@
Figure 7: The path followed by the UAV in the receding horizon jamming implementation, along with the paths of the client, the eavesdropper, and the beampatterns. The UAV initially moves south to get a better angle separation between the client and the eavesdropper, which allows the beampattern (9) to increase. Subsequently, it moves between the client and the eavesdropper and follows them, achieving maximal angle separation while minimizing path losses.

### v2

@@ Abstract
Efficient mobile jamming against eavesdroppers in wireless networks necessitates accurate coordination between mobility and antenna beamforming. We study the coordinated beamforming and control problem for a UAV that carries two omnidirectional antennas, and which uses them to jam an eavesdropper while leaving a friendly client unaffected. The UAV can shape its jamming beampattern by controlling its position, its antennas’ orientation, and the relative phasing for each antenna. We derive a closed-form expression for the antennas’ phases that guarantees zero jamming impact on the client. In addition, we determine the antennas’ orientation and the UAV’s position that maximizes jamming impact on the eavesdropper through an optimal control problem, optimizing the orientation pointwise and the position through the UAV’s control input. Simulations show how this coordinated beamforming and control scheme enables directional GPS denial while guaranteeing zero interference towards a friendly direction.
@@ I Introduction
Two factors largely determine the effectiveness of a jamming signal: the distance to the target and the shape of the antenna’s radiation pattern. Because signal strength decays with distance, staying close to the target ensures sufficient interference power. At the same time, the antenna’s beampattern, i.e., the way its transmitted energy is distributed over different directions, dictates where that power is delivered. A well-shaped beam can focus interference on an eavesdropper at a known direction, whereas a poorly shaped one may also radiate energy toward friendly receivers. Figure 1 illustrates this fact, showing how beamshaping can direct interference selectively at an adversary while avoiding spread towards unintended directions.
Beamshaping for directional jamming has been studied extensively in wireless communication and antenna theory; however, its combination with dynamics is relatively limited. For example, [7] examined beampattern design for static jamming scenarios, while related studies [8, 9, 10, 11, 12] explored different aspects of antenna configuration and beamshaping, all without considering the case where the antenna carrier is a mobile UAV. A related strand incorporates UAV mobility by optimizing trajectories [13, 14, 15, 16], but typically in slot-based formulations that ignore vehicle dynamics and actuator limits, and can generate beampatterns that radiate interference toward friendly users. These approaches highlight the value of beamshaping and UAV mobility, but leave open the question of how to integrate beamforming with dynamics, control, and nulling constraints in a unified framework.
@@
Research in control systems, in contrast, has emphasized the role of mobility in performing or defending against jamming attacks; however, the underlying communication model is often simplified and separated from mobility. For example, [17, 18, 19, 20, 21] presented strategies for planning or agent placement to perform jamming, but assume disk models [^1] The deterministic disk model assumes that jamming is effective within a fixed radius from the receiver, and is otherwise not effective., single omnidirectional antennas, or fixed beampatterns. These provide useful guidelines for the geometry of jammer effectiveness, but restrict capabilities to cases where the target is far from friendly clients or in line of sight; otherwise, jamming the target also jams the friendly client. Other studies have modeled jamming at the control system level [22, 23, 24, 25, 26, 27, 28, 29], but do not account for the communication model that relates spatial geometry to jamming impact. This gap motivates our study of coordinated beamforming and control for directional jamming and nulling.
We consider a unified framework that integrates antenna beamforming with dynamics to achieve directional jamming with guaranteed nulls toward friendly clients. Specifically, we study a jammer UAV that carries two omnidirectional antennas, and which controls their phases and rotational orientation, as well as its position to i) ensure zero jamming power reaches a friendly client, and to ii) maximize jamming power radiated towards an eavesdropper. We assume the antenna spacing is fixed and within one-half of the carrier wavelength, often referred to as a two-element uniform linear array. This configuration allows smooth beampatterns that remain robust to variations in the eavesdropper’s direction (Figure 1).
@@
Notation: $\mathds{1}_{x\in A}(x)$ is the indicator function that is equal to $1$ when $x\in A$, and 0 otherwise. $\left\lVert\cdot\right\rVert$ denotes the Euclidean norm, and $\left\lVert\cdot\right\rVert_{\infty}$ the infinity norm. The function $\textrm{atan2}(y,x):\mathbb{R}\times\mathbb{R}\rightarrow[0,2\pi)$ denotes the two-argument arctangent. We use $I$ to denote an identity matrix of appropriate order, and $\textrm{diag}(c_{1},\ldots,c_{n})$ to denote a diagonal matrix with entries $c_{1},\ldots,c_{n}$. For a vector $x$, $x_{i}$ denotes its $i$-th entry.
@@ II Problem Formulation
We consider a spatial setting where a client wants to receive a wireless communication signal with carrier wavelength $\lambda>0$, while an eavesdropper aims to intercept the same signal. We denote the position of the client in this setting as $p_{c}=[x_{c}~y_{c}]^{\textrm{T}}\in\mathbb{R}^{2}$, and the position of the eavesdropper as $p_{e}=[x_{e}~y_{e}]^{\textrm{T}}\in\mathbb{R}^{2}$.
@@ II-A Spatial Setup for Coordinated Jamming
To prevent the eavesdropper from intercepting the communication, a UAV carries two omnidirectional antennas in order to transmit electromagnetic interference, i.e., to jam the eavesdropper. However, for the jamming to be effective, the UAV must appropriately control the antenna phases of its interference signals as well as its position. Specifically, it must control them to ensure that the interference i) arrives as powerfully as possible at the eavesdropper, hence maximizing jamming impact; and ii) cancels out at the client, hence leaving the client completely unaffected.
We set the distance between the two antennas to be fixed and equal to $0<D\leq\frac{\lambda}{2}$, so that they form a basic two-element antenna array. Given this, if we denote the position of the UAV as $p_{g}=[x_{g}~y_{g}]^{\textrm{T}}\in\mathbb{R}^{2}$, then we can express the positions $p_{1},p_{2}\in\mathbb{R}^{2}$ of each of the antennas as
@@
where $\theta_{g}$ denotes the orientation of the antenna array axis relative to the reference frame at $p_{g}$, which we assume the UAV can mechanically adjust. We provide a geometric illustration of this setup in Figure 2.
@@
Instead of employing a UAV carrying two omnidirectional antennas, one could also employ two UAVs carrying one antenna each. However, this case is more challenging for precise control, antenna array calibration, and beamshape design.
Figure 2: A UAV positioned at $p_{g}=[x_{g}~y_{g}]^{\textrm{T}}$ carries two omnidirectional antennas. The objective is to control its position $p_{g}$, the angle of the two antennas $\theta_{g}$, as well as the antennas’ phases in order to maximally jam the eavesdropper positioned at $p_{e}=[x_{e}~y_{e}]^{\textrm{T}}$. At the same time, a client positioned at $p_{c}=[x_{c}~y_{c}]^{\textrm{T}}$ should remain unaffected.
@@ II-B Jamming Beampatterns and Power
The jamming beampattern generated by the phase selection for the two antennas should ideally cause the two transmitted signals to combine constructively at the eavesdropper $p_{e}$ to maximize jamming effectiveness, and destructively at the client $p_{c}$ to leave it unaffected. At any point $p\in\mathbb{R}^{2}$, this beampattern takes the expression
@@
where $\phi_{1},~\phi_{2}\in\mathbb{R}$ are the beamsteering phases applied at the respective antennas, $d_{1}(p)=\left\lVert p-p_{1}\right\rVert,~d_{2}(p)=\left\lVert p-p_{2}\right\rVert$ are the distances of each of the antennas from $p$, and $k=\frac{2\pi}{\lambda}$.
@@ II-C Coordinated Beamforming and Control for Maximal Jamming
In the case of an immobile UAV, the jamming power at the eavesdropper’s location $p_{e}$ (as given by (3)) can be maximized using standard optimization tools by statically selecting the phases $\phi_{1}$, $\phi_{2}$, and the angle $\theta_{g}$. Here we consider a mobile UAV, where jamming power can be further enhanced by dynamically controlling the UAV position $p_{g}=p_{g}(t)$ over time $t\geq 0$. In this setting, the jamming power (3) becomes time-varying:
@@
and so do the values of the phases $\phi_{1}(t),\phi_{2}(t)$ and the angle $\theta_{g}(t)$, allowing for greater jamming flexibility. Controlling the UAV position $p_{g}(t)$ while concurrently selecting the antennas’ transmission parameters $\theta_{g}(t),\phi_{1}(t),\phi_{2}(t)$ is the purpose of the remainder of the paper.
@@
$\quad\qquad\left\lVert u_{g}(t)\right\rVert_{\infty}\leq\bar{u},~\forall t\in[0,t_{f}],$
@@
While the free-space path losses term $L$ indicate that smaller distances lead to higher jamming impact, it is evident from (4) that such impact also depends on the beamforming function $B$. This means that the array formed by the UAV’s antennas should reposition not only to get closer to the eavesdropper, but also arrive from the proper direction and with the appropriate orientation $\theta_{g}$. For this reason, the position $p_{g}$ of the UAV is not included explicitly in (5) but, instead, its desired values are informed through the jamming impact power $P$.
@@ III-A Phase Control for Nulling at the Client
guarantees the client remains unaffected by the jamming. Subsequently, we may choose the phase $\phi_{1}(t)$ arbitrarily: since it represents just an offset on the phase of the beamformed signal on the eavesdropper, it has no effect on the magnitude of the beampattern (1). However, note that for the narrow band signal case considered here, we can adjust $\phi_{1}(t)$ to offset for Doppler shift $f_{D}$ on the jammer according to
@@ III-C Position Control for Maximal Jamming Impact on the Eavesdropper
Having chosen the jamming phases $\phi_{1}(t),~\phi_{2}(t)$ and the array orientation $\theta_{g}(t)$, the final step is to select the control input $u_{g}(t)$ that repositions the UAV for maximal jamming impact. To this end, plugging the expression of the optimal antenna orientation (10) in (9), we obtain the optimal beampattern value:
@@
$\begin{split}{\xi}_{p}(t_{f})&=-a_{f}\gamma(t_{f})\Bigg{(}\frac{B_{p}(p_{g}(t_{f}))\cdot\mathds{1}_{|\mu(t_{f})|<\frac{\pi}{2kD}}}{B^{\star}(\theta_{e}(t_{f});p_{g}(t_{f}))}\\ &\qquad\qquad\qquad\quad{+}\frac{L_{p}(p_{g}(t_{f}))}{L(\left\lVert p-p_{g}(t_{f})\right\rVert)}\Bigg{)},\\ {\xi}_{v}(t_{f})&=Q_{f}v_{g}(t_{f}).\end{split}$
@@ IV-A Directional Jamming Against a Static Eavesdropper
To jam the eavesdropper and prevent it from receiving the wireless signal, we deploy a jammer UAV that carries two omnidirectional antennas separated by $D=\frac{\lambda}{2}=9.52~\textrm{cm}$ and with power $P_{0}=600~\textrm{mW}$ each, as shown in Figure 2. This UAV is initially positioned at $p_{g}(0)=[0~0]^{\textrm{T}}~\textrm{m}$ and with velocity $v_{g}(0)=[0~0]^{\textrm{T}}~\textrm{m/s}$. This position, however, is suboptimal: the UAV is far from the eavesdropper, and both the client and the eavesdropper lie in the same direction in the far field. As a result, the UAV cannot effectively jam the eavesdropper without also interfering with the client. Therefore, to reposition the UAV for maximum jamming effectiveness, we solve the optimal control problem (14) to determine its control input $u_{g}$, and we select the communication parameters $\phi_{1},~\phi_{2},~\theta_{g}$ according to (6), (7) and (10). The actuation threshold for the UAV is $\bar{u}=2~\textrm{m/s}^{2}$.
@@
and, accordingly, $\sigma^{\prime}(x)=\mathds{1}_{x\in[-100,-70]}(x)$. This choice of $\sigma$ motivates the optimal $u_{g}^{\star}$ to generate trajectories that degrade or deny service at the eavesdropper. At the same time, it avoids giving excessive rewards for UAV paths that pass too close to the eavesdropper and yield jamming powers above $70~\textrm{dBm}$ (as the DoS outcome remains unchanged). Moreover, it preserves numerical stability by zeroing the flow equations in Theorem 2 whenever either $B^{\star}$ or $\left\lVert p-p_{g}\right\rVert$ is close to zero. Note, to improve numerical stability, we use the smooth approximation $\sigma^{\prime}(x)\approx\frac{1}{1+e^{-200(x+100)}}-\frac{1}{1+e^{-200(x+70)}}$.
@@
Figure 7: The path followed by the UAV in the receding horizon jamming implementation, with the client denoted in steel blue and the eavesdropper in dark red. The UAV initially moves south to get a better angle separation between the client and the eavesdropper, which allows the beampattern (9) to increase. Subsequently, it moves between the client and the eavesdropper and follows them, achieving maximal angle separation while minimizing path losses.